### v1

# PLANesT-3D: A new annotated dataset for segmentation of 3D plant point clouds

Abstract
Creation of new annotated public datasets is crucial in helping advances in 3D computer vision and machine learning meet their full potential for automatic interpretation of 3D plant models. In this paper, we introduce PLANesT-3D; a new annotated dataset of 3D color point clouds of plants. PLANesT-3D is composed of 34 point cloud models representing 34 real plants from three different plant species: Capsicum annuum, Rosa kordana, and Ribes rubrum. Both semantic labels in terms of "leaf" and "stem", and organ instance labels were manually annotated for the full point clouds. As an additional contribution, SP-LSCnet, a novel semantic segmentation method that is a combination of unsupervised superpoint extraction and a 3D point-based deep learning approach is introduced and evaluated on the new dataset. Two existing deep neural network architectures, PointNet++ and RoseSegNet were also tested on the point clouds of PLANesT-3D for semantic segmentation.
Keywords 3D point cloud  $\cdot$ 3D plant modeling  $\cdot$ 3D plant dataset  $\cdot$ 3D semantic segmentation
1 Introduction
Cultivation of horticultural crops is one of the most labor-intensive economic activities, and is central to global food security [1, 2]. The accuracy and efficiency of horticultural practices, such as pruning, weed control, disease management, and harvesting directly influence the yield and quality of the crops [3]. The increasing global demand for high-quality crops, hence the demand for intensive labor motivates development of technologies for automation in horticulture [4, 5]. Automated horticultural practices are expected to adapt to plant traits, which are diverse across plant species [6, 7]. Automation of trait measurements is also key to high throughput plant phenotyping, a fundamental process in plant research, breeding, and crop management [8, 9, 10]. Modeling of traits within a population, and capturing the structure of single plants or plant communities for particular operations have become major applications of computer vision [11, 12, 13, 14].
3D computer vision is particularly relevant in many automated horticultural operations. Robotic systems designed to perform pruning, weeding, and harvesting require correct identification of plant parts, and accurate estimation of the location, size, shape, and orientation of the structures in 3D space [15, 16, 17, 18]. Advances in 3D sensing technologies and machine learning fuel the development of algorithms for automatic segmentation of plants into their structural parts as well as organ-level trait estimation [19, 20].
Despite the considerable progress in 3D geometry modeling and analysis of plants in the last decades [21, 22], creation of benchmark datasets, a fundamental component for advancement of relevant research, is still lagging behind compared to other domains [23, 24]. Annotated public datasets of plant models are of particular importance both for comparison of competing algorithms and also for providing training data to machine learning processes [13].
Table 1: Public datasets containing full 3D models of plants in comparison with the PLANesT-3D dataset introduced in this work.
dataset Properties
# models Species Modality Color Labels
Pheno4D 126 Tomato Laser No Semantic &
[25] Maize Instance
ROSE-X 11 Rose X-rays No Semantic
[26]
Plant3D 714 Tomato Laser No
 [27] Tobacco Growth
[28] Sorghum conditions
[29] Arabidopsis
Soybean-MVS 102 Soybean MVS Yes Semantic &
[30] Instance
PLANesT-3D 34 Pepper Yes Semantic
Rose SfM-MVS &
Ribes Instance
In the last decade, open datasets of annotated 2D plant images have been proliferated to assist evaluation of performance and development of machine learning models for plant phenotyping [31]. However, open datasets containing annotated and complete 3D plant models are rare. Table 1 gives a list of such existing datasets. Plant3D [27, 28, 29] contains a total of 714 3D laser scans of Tomato, Tobacco, Sorghum, and Arabidopsis plants obtained within 20-30 days of development. Information on the growth conditions of the plants are also provided. The 3D models do not contain color information. [32] manually labeled 54 models of the Plant3D data into stem and lamina points, and evaluated a range of machine learning approaches on this subset. [33] annotated 546 of the plants of the Plant3D dataset with semantic and instance labels. The labels are available upon request from [33]. ROSE-X dataset [26] contains 3D models of 11 rosebush plants acquired through X-ray computer tomography. The voxels in the volumetric models were labeled into three semantic categories: "Leaf", "Stem", and "Flower". The point cloud versions of the volumetric models are also available. Pheno4D [25] is a dataset of 3D point clouds of 7 maize plants and 7 tomato plants. The plants were scanned with a laser scanner at different growth stages resulting in 244 point clouds. 126 of them were manually annotated with semantic and instance labels. The Soybean-MVS dataset [30] is fundamentally different from the rest in terms of data acquisition modality. The plants were captured with an RGB camera in a controlled setup and their corresponding point clouds were created through multi-view stereo. A total of 102 point cloud models of five different soybean varieties were reconstructed at 13 stages of the whole growth period.
These 3D plant datasets have been often used for assessing off-the-shelf machine learning tools for semantic and instance segmentation [32, 25, 34, 30]. The datasets also enabled training and evaluation of new approaches specifically developed for plant data [35, 36, 37, 38], including deep learning architectures such as PSegNet [33], PlantNet [39], RoseSegNet [40], and FF-Net [41].
Although smaller in terms of the number of models as compared to Pheno4D, Plant3D, and Soybean-MVS (Table 1), the new dataset (PLANesT-3D  [^1] The PLANesT-3D dataset is publicly available at https://www.dropbox.com/s/8gl0r0zo0j6kmdo/PLANesT-3D.rar?dl=0) introduced in this work contributes to the diversity of publicly available annotated 3D plant models in terms of acquisition modality, number of distinct plants, and inclusion of other species. The data is acquired with an RGB camera and reconstructed through structure from motion (SfM) and multi-view steoroscopy (MVS). Together with the Soybean-MVS dataset, the PLANesT-3D dataset will balance the dominant acquisition modality, which is laser scanning. Reconstruction with SfM and MVS does not require expensive equipment or elaborate acquisition setups, while yielding both geometric and textural information.
While Pheno4D and Soybean-MVS datasets are extremely valuable since they track the development of individual plants across time, the number of distinct plants are 14 for Pheno4D and 5 for Soybean-MVS. The PLANesT-3D dataset includes models of 34 distinct plants belonging to three species: namely Capsicum annuum (pepper), Rosa kordana (rose), and Ribes rubrum (ribes). Also, point clouds from two species (pepper and ribes), 3D models of which were not available before, are added to the public domain. The diversity of training and evaluation data in terms of acquisition modality, noise levels, plant species, plant instances, plant architecture, organ geometry, etc. plays an important role in assessing the robustness as well as the generalization ability of machine learning approaches.
The objective of this publication is twofold: First is to introduce PLANesT-3D, a new dataset containing complete 3D color point clouds of plants together with their semantic and instance labels. Second is to present a new segmentation approach, the SP-LSCnet, that combines an unsupervised clustering scheme and an adaptive network for point cloud classification. We also provide semantic segmentation results on the dataset with two point-based deep learning architectures, PointNet++ and RoseSegNet. The main contributions of this work can be summarized as follows:
  1. 1.
    We introduce PLANesT-3D, a new dataset containing RGB point clouds representing 34 real plants from three species.
  2. 2.
    We propose SP-LSCnet as a new semantic segmentation approach where an unsupervised clustering method based on t-distributed stochastic embedding (t-SNE) is combined with a point cloud classifier network. The classifier network employs two adaptive modules to adjust local region organization for feature extraction.
  3. 3.
    We tested RoseSegNet, a semantic segmentation algorithm we had previously developed for plant data [40] on the new dataset, and demonstrated that it is effective without requiring hyperparameter re-adjustment.
The rest of this article is organized as follows. In Sections 2.1 and 2.2, we describe the data acquisition, reconstruction, and pre-processing steps leading to the creation of 3D point clouds. We give detailed information about the PLANesT-3D dataset in Section 2.3. The segmentation methods evaluated in this work are described in Section 2.4. Our new segmentation approach, SP-LSCnet, is introduced in Section 2.4.1. The performances of the semantic segmentation methods on the PLANesT-3D dataset are reported in Section 3. A discussion of the work is provided in Section 4, followed by our conclusion in Section 5.
Figure 1: Estimated camera poses for 231 images of a pepper plant (a), 240 images of a rose plant (b), and 177 images of two ribes plants (c). Sample images for pepper (d), rose (e), and ribes (f) are also provided.
2 Materials and Methods
2.1 Data acquisition
In order to obtain full 3D plant models, 2D color images of 34 plants were acquired using a handheld DSLR camera (EOS 6D MARK II, Canon, Japan). The acquisition was performed on 10 potted pepper plants (Capsicum annuum), 10 potted rose plants (Rosa kordana), and 14 ribes plants (Ribes rubrum), planted directly to the soil in a greenhouse. Multiple color images of resolution $6240\times 4160$ were captured manually around each plant via positioning the camera so as to cover as much plant surface as possible with a high degree of overlap. Three examples of the sets of camera locations and orientations are illustrated in Fig. 1.
Table 2 gives the number of images used to reconstruct the 3D plant point clouds. Three pairs of Ribes were very closely planted (Ribes 02 & 12, Ribes 04 & 13, Ribes 05 & 14). A single point cloud was reconstructed for each pair using the set of images where the pair appeared together. After reconstruction, noise removal and pose normalization, the two plants were manually segmented into separate point clouds of two individual plants.
Table 2: Number of images used to reconstruct the 3D plant point clouds.
Plant ID # images Plant ID # images Plant ID # images
Pepper 01 261 Rose 01 210 Ribes 01 243
Pepper 02 320 Rose 02 182 Ribes 02 & 12 177
Pepper 03 324 Rose 03 203 Ribes 03 294
Pepper 04 282 Rose 04 205 Ribes 04 & 13 235
Pepper 05 280 Rose 05 240 Ribes 05 & 14 169
Pepper 06 231 Rose 06 211 Ribes 06 150
Pepper 07 294 Rose 07 258 Ribes 07 225
Pepper 08 350 Rose 08 406 Ribes 08 222
Pepper 09 199 Rose 09 377 Ribes 09 191
Pepper 10 226 Rose 10 386 Ribes 10 310
Ribes 11 240
2.2 Reconstruction and pre-processing of 3D point clouds
For reconstruction of 3D color point clouds from 2D color images, Agisoft Metashape Professional (Agisoft LLC, St. Petersburg, Russia) was employed. Agisoft Metashape is a software that performs photogrammetric processing of digital images and generates 3D models through principles of structure from motion (SFM) and multi-view reconstruction [42]. Examples of the raw point clouds produced by Metashape Professional are given in Fig. 2. The software also provides confidence maps for the 3D points. The confidence value for a point corresponds to the number of images that "see" and contribute to reconstruct the point.
As can be observed from Fig. 2, the raw point clouds include structures from the background and are noisy. The correct scale of the scene is lost. Automatic segmentation of plants from each other and from the background is in itself a research problem. However, we aimed to provide a dataset of 3D models that correspond to isolated plants for evaluation of algorithms that focus on the analysis of single plants.
Figure 2: Raw point clouds reconstructed via Agisoft Metashape Professional of a rose plant (a) and a ribes plant (c). Corresponding clouds including only 3D points of the target plants are given in (b) for rose, and in (d) for ribes.
In order to recover the correct scale of the point cloud, remove background structures and noise, and segment the plant points from the scene, we devised a simple, semi-automatic procedure. We denote the raw point cloud as $\mathcal{P}_{R}$, where a point $p\in\mathcal{P}_{R}$ is represented by its coordinates $p=(x,y,z)$ defined in the coordinate frame returned by Metashape. The steps of the procedure are detailed as follows:
[…]
    $\hat{D}_{R}=\frac{1}{N_{r}(N_{r}-1)}\sum_{r\neq s}{\sum_{s}{\hat{d}_{rs}}}\;.$ (1)
    Let $D_{R}$ be calculated using the true metric distances (in cm units) between the landmark points $d_{rs}$ measured manually, or read through the ruler. The scale factor is then determined as
    $S=\frac{D_{R}}{\hat{D}_{R}}\;.$ (2)
    The point cloud is scaled such that the average of the reference distances on the point cloud is equal to the measured distances on the real objects, i.e. the point coordinates were updated as $p^{\prime}=S\cdot p$ for all $p\in\mathcal{P}_{R}$. The scaled point set is denoted as $\mathcal{P}^{\prime}_{R}$.
  • Rotating the point cloud to a normalized pose: We aimed to rotate the reference frame such that the XY-plane corresponded to the ground or the plane holding the plant pot. In order to estimate the parameters of this plane, we used M-estimator SAmple Consensus (MSAC) algorithm given in [43], which is a variant of RANdom SAmple Consensus (RANSAC) algorithm. We manually picked a point, $p^{\prime}_{base}$, close to the plant base. Using the estimated normal vector $\mathbf{n}=[n_{x},n_{y},n_{z}]$ of the plane in the scene and the point $p^{\prime}_{base}$ at the plant base, we translated and rotated the point cloud $\mathcal{P}^{\prime}_{R}$ into a new reference frame using
    $\hat{p}=\mathbf{R}(p^{\prime}-p^{\prime}_{base})$ (3)
    for all $p^{\prime}\in\mathcal{P}^{\prime}_{R}$. The point cloud rotated to this normalized pose is denoted as $\hat{\mathcal{P}}_{R}$. In Eq. 3, $\mathbf{R}$ corresponds to the rotation matrix which transforms the normal $\mathbf{n}$ to $[0,0,1]$ in the new reference frame. The origin of this reference frame corresponds to the hand-picked point at the plant base. The XY plane coincides with the plane detected by the MSAC algorithm. The positive Z-direction is oriented from the plant base towards the shoot.
  • Extracting plant points: Once the origin of the reference coordinate frame is established at the plant base, and the positive Z-direction points toward the plant shoot, we removed all points with a negative $z$ coordinate, assuming all plant points remain above the plant base. We verified this assumption for all the models in the dataset. We applied connected component analysis to the rest of the point cloud and retained the points belonging to the largest connected component as the final plant point set $\mathcal{P}\subset\hat{\mathcal{P}}_{R}$. Examples of the outputs of this semi-automatic process is given in Fig. 2-(b) for a rose plant and in Fig. 2-(d) for a ribes plant.
Figure 3: Three models from the PLANesT-3D dataset rendered with plant color (a), confidence map (b), semantic labels (c), and instance labels (d).
2.3 The PLANesT-3D dataset
The PLANesT-3D dataset consists of 34 plant point clouds, where the plant points are isolated from the background as described in the previous section. Accompanying the locations and color information of the points in each model, three scalar fields are provided: The confidence map as delivered by Metashape Professional, semantic labels, and instance labels. Examples of point clouds from the PLANesT-3D dataset and their corresponding confidence maps and labels are depicted in Fig. 3.
Both semantic labels and instance labels are obtained through manually labeling each point using CloudCompare point cloud processing software [44]. Semantic labels for PLANesT-3D correspond to "leaf" and "stem" classes. Points on the leaf blades were labeled as "leaf", main stem, branches, and petioles were included in the "stem" class. Instance labels are identity numbers representing individual leaflets within a plant model. All points labeled as "stem" were assigned a single instance label.
Table 3: Properties of the point clouds in PLANesT-3D
Height (cm) # points "leaf" (%) "stem" (%) # leaflets
Pepper 01 53.63 2312959 70.45 29.56 33
Pepper 02 63.39 5531852 76.80 23.20 48
Pepper 03 60.65 3482959 72.43 27.57 45
Pepper 04 51.51 2425889 68.28 31.72 40
Pepper 05 48.95 2720311 69.52 30.48 35
Pepper 06 32.94 1846694 65.41 34.59 21
Pepper 07 49.71 1806010 69.11 30.89 27
Pepper 08 48.45 2311602 69.46 30.54 27
Pepper 09 53.32 3440529 74.10 25.90 38
Pepper 10 64.95 2416062 69.91 30.09 36
Rose 01 29.96 3452545 80.15 19.85 59
Rose 02 33.55 3601210 80.55 19.45 68
Rose 03 32.96 2580315 75.97 24.03 43
Rose 04 27.36 4091952 82.37 17.63 61
Rose 05 38.01 2867159 78.29 21.71 65
Rose 06 25.98 2540487 81.62 18.38 27
Rose 07 27.88 2999376 82.24 17.76 43
Rose 08 46.80 3616281 74.12 25.88 100
Rose 09 50.16 2119400 77.45 22.55 56
Rose 10 37.65 3296993 77.65 22.35 85
Ribes 01 76.92 2078729 88.33 11.67 64
Ribes 02 33.69 1299910 89.67 10.33 34
Ribes 03 74.25 1053341 84.32 15.68 33
Ribes 04 34.64 1190366 85.32 14.68 32
Ribes 05 104.85 1506387 78.09 21.91 105
Ribes 06 31.55 3175798 81.49 18.51 63
Ribes 07 93.66 3553784 85.66 14.34 101
Ribes 08 79.73 600792 67.04 32.96 41
Ribes 09 43.53 679885 68.20 31.80 34
Ribes 10 62.12 1256866 82.29 17.71 54
Ribes 11 33.32 1194023 61.12 38.88 99
Ribes 12 52.52 3147426 87.08 12.92 70
Ribes 13 46.18 3258674 79.56 20.44 92
Ribes 14 56.13 644626 83.33 16.67 21
Table 3 gives five properties of each of the 34 plant point clouds in the dataset: 1) Plant height, 2) Total number of points, 3) Percentage of "leaf" points, 4) Percentage of "stem" points, and 5) Number of individual leaflets. Plant height is in the range of 25 to 105cm. The number of points exceed 1 million for most plants. Number of individual leaflets varies between 25 to 105, indicating the variation of complexity in plant structure.
2.4 Methods for semantic segmentation
Harandi et al. [22] provided a comprehensive review of typical steps involved in the processing and analysis of 3D representations of plants, including point clouds. Models in PLANesT-3D can be used to evaluate a diverse set of tools developed for plant analysis and phenotyping. In this study, we focus on the application of segmentation of 3D point clouds of plants into their semantic units.
The objective of semantic segmentation is to assign each point $p$ in the point set $\mathcal{P}$ one of the semantic labels. The semantic categories in the PLANesT-3D dataset correspond to "leaf" and "stem", where "leaf" points are on leaf blades, and "stem" points belong to the main stem, branches, and petioles.
We developed a novel semantic segmentation method that is a combination of a superpoint extraction scheme [38] and an adaptive 3D object classification network [45]. We abbreviate this method as SP-LSCnet, where SP stands for superpoint, while LSCnet is a leaf-stem classification network. In order to provide reference performance results for further research, we also tested PointNet++ [46] and RoseSegNet [40] on the new PLANesT-3D dataset.
2.4.1 SP-LSCnet
SP-LSCnet consists of two stages: Superpoint extraction and leaf-stem classification. In superpoint extraction stage, the point cloud is over-segmented into superpoints through an unsupervised process that operates on 2D points embedded by t-SNE. Then, each superpoint generated in the first stage is classified through a deep neural network equipped with attention-based modules that adaptively determine receptive fields for feature extraction. The flowchart of the method is depicted in Fig. 4, and the details of the two stages are given below:
Figure 4: Flowchart of SP-LSCnet. The input point cloud is first passed through the superpoint extraction procedure. This procedure involves embedding of the 3D points into 2D and partitioning the 2D points into superpoints. Once the superpoint assignments are mapped to 3D domain, each superpoint is passed independently through a leaf-stem classification network. The 3D points in a superpoint are all labeled with the global class of the superpoint predicted by the network.
Superpoint extraction: For accomplishing this task, the superpoint extraction scheme proposed in [38] is selected. The plant point cloud is first downsampled with voxel grid average filtering. t-SNE, as introduced in [47], is used to embed the downsampled 3D point cloud into 2D space. Superpoint extraction is then performed exclusively in 2D domain. t-SNE inherently provides clusters that are formed based on the implicit manifolds to which the 3D points belong.
The first step of the superpoint extraction scheme corresponds to the identification of these clusters by Euclidean clustering in 2D. Line-like structures are detected through examining the local geometry of the points. An iterative procedure that involves spectral clustering and solidity computation is employed to segment non-convex clusters into convex regions. Once all 2D points are assigned to their corresponding superpoints, the assignments are carried to their 3D counterparts, and propagated to the high-resolution point cloud through nearest neighbor interpolation.
In this work, for the PLANesT-3D dataset, the grid size for voxel grid average filtering is selected as 0.12cm, the $Perplexity$ for t-sNE embedding is set to 30. The distance threshold used in Euclidean clustering in 2D space is chosen as 1. The other parameters were kept in their default settings as given in [38].
Figure 5: The integration of CSM and RUM to PointNet++ framework [45]: a) The classification network. $r^{L}$ for $L=1$ (first layer) is set to be 0.2. $r^{L}$ for $L=2$ (second layer) is 0.4. b) Adaptive local region inference with CSM and RUM.
Leaf-stem classification network: Each superpoint generated by the previous step is considered to be a separate point cloud. A point-based deep neural network, introduced by [45], is utilized to classify the superpoints into leaf and stem classes (Fig. 4). The network is formed through integrating two modules to the PointNet++ architecture for object classification (Fig. 5). The modules are based on attention mechanisms to analyze point interactions within and between point neighborhoods, and are named as Center Shift Module (CSM) and Radius Update Module (RUM) [45]. Their task is to update adaptively the centers and radii of the spherical regions on which PointNet++ encodes local features. This operation is referred to as adaptive local region inference [45].
The classification network architecture of PointNet++ [46], shown in Fig. 5(a), includes three set abstraction (SA) layers. The input point cloud with $N=1024$ points is reduced to $N_{1}=512$, and $N_{2}=128$ representative points during SA layer-1 and SA layer-2, respectively. At each SA layer $L$, the representative points $c_{j}^{L}$ are selected using farthest point sampling (FPS) algorithm. $N_{L}$ spherical local regions with radius $r^{L}$ is defined around the center points $c_{j}^{L}$. Within each region, $K$ points are randomly selected and the features of $K$ points are mapped to higher dimensions with MLPs. The abstracted features of each local region are calculated by taking the maximum among the feature channels of $K$ grouped points.
The global descriptors representing the input point cloud are extracted at the last layer, where a single region enclosing the entire object is encoded. The global descriptors are passed through fully connected layers to compute the category scores. In this study, the classification network calculates two category scores for each input point cloud: One for the "leaf" category, and one for the "stem" category.
The adaptive local inference modules, CSM and RUM can be integrated to either or both of the first two layers as shown in Fig. 5(b). If CSM is "ON" at a particular layer, for each representative point $c_{j}^{L}$, CSM calculates the shift amount $\Delta c_{j}^{L}$, through attention-based subnetworks. The representative point is updated as:
$\hat{c}_{j}^{L}=c_{j}^{L}+\Delta c_{j}^{L}\;.$ (4)
If RUM is "ON", it computes the amount of radius update $\Delta r_{j}^{L}$ for each region $j$. The new radius then becomes:
$\hat{r}_{j}^{L}=r^{L}+\Delta r_{j}^{L}\;.$ (5)
The grouping and feature encoding by MLPs of $K$ points are then performed using these new region definitions. These modules allow the receptive fields to be adaptively shifted and resized through examining local and global point interactions through attention. The network thus adjusts the receptive fields and encodes the 3D points in them in accordance with the main task of the network.
Various alternatives of CSM and RUM were suggested by [45]. As a result of experimental trials, the CSM-II (sub) and RUM-II (cum) versions were selected for the PLANesT-3D dataset. Please, refer to [45] for detailed descriptions of of the modules CSM-II (sub) and RUM-II (cum).
2.4.2 PointNet++
We adopted the PointNet++ architecture [46] for semantic segmentation of large point clouds. The data preparation for PointNet++ involves two operations: Subsampling and partitioning the big point set into subsets enclosed by blocks.
The subsampling is performed using voxel-grid filtering, where the edge length of the voxel cube is set as $1mm$. Each voxel is represented with a single point which is the average of the original points occupying the voxel. The sampled point is assigned the label of the closest point in the original cloud. This subsampling operation produces a point cloud with reduced size and homogeneous point density.
3D point-based deep neural networks requires input clouds with a fixed number of points. One way is to subsample the full plant point cloud to have the required number of points. However, such strategy will result in significant loss of geometric information, especially for small structures. Instead, we selected to partition the plant point cloud into subsets of points, each of which is then processed by the network individually. We followed a similar procedure of partitioning the point clouds into blocks as in [34, 40].
The horizontal region, i.e. the extent of the $XY$ plane, encompassing the point cloud is partitioned into squares of edge length of $10cm$. All points with $x$ and $y$ coordinates falling into a square region is considered belonging to the point subset within a block. The points in the block are subsampled to obtain a cloud of fixed number of points, as $N=8192$. The partitioning operation is conducted twice for each plant point cloud with offset values 00 and $5cm$.
We used the default PointNet++ architecture for semantic segmentation as given in [46]. PointNet++ segmentation network is composed of five grouping/abstraction layers. The radius parameter, which defines the size of local regions for grouping and abstraction, is set for each of the first four layers as $0.5cm$, $1cm$, $2cm$, and $4cm$, respectively. During training, batch size was selected as 16. The learning rate was updated as 0.005. Other hyperparameters of the PointNet++ were kept at default values.
2.4.3 RoseSegNet
RoseSegNet was developed for the specific application of semantic segmentation of rosebush models. It involves attention-based modules that encode point interactions within and between local regions. The details of the RoseSegNet architecture can be found in [40].
As with the case of PointNet++, we partition the plant point cloud into blocks and the point subset in each block is separately processed by RoseSegNet. The block partitioning process is the same as with PointNet++.
All the hyperparameters of the RoseSegNet are kept at their default values as given in [40], where the parameter search was performed on the ROSE-x dataset [26].
3 Results
In this section, we provide semantic segmentation results on the PLANesT-3D dataset obtained by the three methods described in Section 2.4. We trained a separate network for each of the three plant species. The hyperparameters of the networks were kept fixed for all the species. 70% of the point clouds were selected for training the networks. The rest is reserved for test. The split for training and test sets for the three species are given in Table 4. The IDs of the plants in the test sets are also provided.
Table 4: The split for training and test sets
Species Pepper Rose Ribes
# plants for training 7 7 10
# plants for test 3 3 4
Plant IDs reserved for test 01, 03, 07 01, 03, 09 03, 10, 11, 14
Table 5: Number of point subsets used for training and test.
Species Pepper Rose Ribes
# training blocks for PointNet++ & RoseSegNet 294 317 665
# test blocks for PointNet++ & RoseSegNet 123 106 205
# training superpoints for SP-LSCnet 1188 933 2761
# test superpoints for SP-LSCnet 646 438 998
Before being processed by the networks, the point clouds were filtered such that points with confidence values less than 6 were discarded. For PointNet++ and RoseSegNet, the point cloud of a plant is partitioned into blocks and each block is separately processes as described in Section 2.4.2. For SP-LSCnet, the point clouds are partitioned into superpoints, and the superpoints are treated as separate objects for the object classification network. Table 5 gives the number of point subsets used for training and test for each plant species.
Table 6: Semantic segmentation results of the three methods on the PLANesT-3D dataset.
PEPPER
PointNet++ RoseSegNet SP-LSCnet
Precision - Stem 96.5 97.0 97.6
Recall - Stem 94.6 96.1 94.5
IoU - Stem 91.5 93.2 92.4
Precision - Leaf 98.3 98.7 98.3
Recall - Leaf 98.9 99.0 99.3
IoU - Leaf 97.2 97.8 97.6
Acc 97.9 98.3 98.1
MIoU 94.3 95.5 95.0
ROSE
PointNet++ RoseSegNet SP-LSCnet
Precision - Stem 92.4 94.8 96.4
Recall - Stem 89.7 91.4 85.6
IoU - Stem 83.5 87.1 83.0
Precision - Leaf 97.6 98.0 97.2
Recall - Leaf 98.3 98.8 99.4
IoU - Leaf 96.0 96.9 96.6
Acc 96.7 97.4 97.1
MIoU 89.8 92.0 89.8
RIBES
PointNet++ RoseSegNet SP-LSCnet
Precision - Stem 95.8 95.9 96.7
Recall - Stem 94.0 95.5 94.0
IoU - Stem 90.3 91.7 91.1
Precision - Leaf 98.9 99.2 98.5
Recall - Leaf 99.2 99.2 99.2
IoU - Leaf 98.1 98.4 97.8
Acc 98.4 98.6 98.3
MIoU 94.2 95.1 94.5
We report the semantic segmentation results of the three methods in Table 6. Precision, Recall, and Intersection over Union (IoU) measures specific to "stem" and "leaf" classes are given in the table, as well as overall accuracy (Acc) and mean IoU (MIoU). In terms of overall accuracy and mean IoU, RoseSegNet yielded the best performance for all three species. For MIoU, RoseSegNet surpassed the performance of PointNet++ by close to 1%. SP-LSCnet is in between, performing slightly better than PointNet++ for the Pepper and Ribes sets, while yielding similar peformance to that of PointNet++ for the Rose set.
Figure 6: Semantic segmentation results by PointNet++, RoseSegNet, SP-LSCnet, on sample pepper (a), rose (b), and, ribes (c) plants. First column corresponds to the ground truth.
The results on rose plant models are lower compared to those obtained on pepper and ribes plants. The recall of the "stem" class is more challenging for the rose plant models, since the petioles in between the leaflets are hard to discern. The misclassifications of petiole points of roses as belonging to leaf blades are observable in Fig. 6 and Fig. 7 where segmentation results are visually depicted for sample plants.
Figure 7: Examples to errors
Other examples to erroneous cases are given in Fig. 7. The errors produced by PointNet++ and RoseSegNet appear as isolated points scattered in particular regions. Although context information heavily influences the features of single points in PointNet++ and RoseSegNet, the classification is performed individually for each point. Misclassifications occur particularly at cluttered regions where leaves and stems are close to each other.
The treatment of each superpoint as a single entity by SP-LSCnet produces smooth results; however, it suffers from two disadvantages: First, the superpoints might not be always homogeneous; some leakage from other classes is inevitable due to the unsupervised nature of superpoint extraction. Second, the correct classification of the superpoints becomes particularly critical. As can be seen in Fig. 7, portions of the stem were classified as leaf points if the single superpoint to which they belong has the geometrical characteristics of leaves as modeled by the network. Likewise, small leaves, where the point resolution is not sufficient for proper sampling of the leaf surface, are classified as stem. Low point resolution due to occlusion (see bottom-left example in Fig. 7) can also cause isolated regions separated into superpoints which are not large enough to possess distinguishing shapes.
All three methods can benefit from post-processing such as label smoothing for PointNet++ and RoseSegNet, and graph-based re-evaluation of superpoint predictions for SP-LSCnet. However, the limitations of these methods in dealing with scarce annotated data, noisy point clouds, complex architecture of plants, and variable point density call for further developments in plant segmentation methods.
4 Discussion
The PLANesT-3D dataset is composed of 3D point clouds of plants reconstructed through structure from motion and multi-view stereo. Capturing 3D geometry through SfM has many advantages such as low cost setup requirements, high resolution, and provision of color information [22]. It can be used both in controlled environments and in the field without extensive training for acquisition and elaborate protocols. A public dataset obtained through this modality as opposed to laser scanning is expected to be instrumental in assessing the robustness of 3D plant processing and phenotyping tools.
As can be observed in Table 3, the size of each model in terms of number of points is high, ranging from 600K to 5.5M points. One question is whether such high resolution is necessary for certain tasks related to plant analysis and management. The number of points can be reduced systematically through downsampling to measure the effect of resolution on the performance of particular 3D shape processing tools [48].
We also used quite large numbers of images of the plants for reconstruction (2). This amount of image capture might not be feasible in cases where manual data acquisition is performed. However, advances in robotics enable fast acquisition with autonomous crop surveying robots equipped with robotic arms and multiple cameras [8, 49, 50], as well as techniques for automatic positioning of the robot arms for optimal data acquisition [51]. Nevertheless, limiting the number of images for reconstruction without a performance degradation in particular tasks would substantially reduce the computational demand.
Another important aspect of the new dataset is the availability of color information. Although not demonstrated in this work, use of color information together with 3D geometry can boost the performance of a segmentation procedure [52]. The PLANesT-3D dataset will be a valuable resource for measuring the contribution of color information for automated plant analysis and phenotyping tasks.
Semantic parts annotated in the PLANesT-3D dataset are limited to "leaf" and "stem" classes; however, a more fine-grained annotation is possible, for example, in terms of "leaf blades", "petioles", "main stem", "nodes", and "internodes" [53]. Leaflets, together with petioles, can be grouped into their corresponding individual leaves. The architecture of each plant can be modeled with a graph representation. We envision undertaking this level of annotation as a future work.
Another possible future work is a study on the transferability of the neural network models learned from a plant species into another species. In the current work, we trained a separate network for each of the three plant species. The extent of generalization ability of network models among plant species is of special interest; since manual annotation of plants for training is highly labor-intensive. PLANesT-3D, with a collection of three different species, allows application and evaluation of algorithms such as small sample learning and, transfer learning.
5 Conclusions
We introduced the PLANesT-3D dataset, which is composed of annotated color point clouds of 34 plants belonging to three different species. We described the acquisition, reconstruction, and labeling processes involved in dataset construction. As a use case of the dataset, we reported semantic segmentation results yielded by PointNet++, RoseSegNet, and a novel method abbreviated as SP-LSCnet. RoseSegNet achieved highest segmentation performance, reaching 95.5 %, 92.0 %, and 95.1 % MIoU, for pepper, rose, and ribes plants, respectively. SP-LSCnet also produced comparable results as 95.0 %, 89.8 %, and 94.5 % MIoU on the three plant species. We illustrated sample cases where the three methods failed to produce correct results. The PLANesT-3D dataset will be instrumental for developing new methods that effectively address these challenges.
Acknowledgements:
The authors acknowledge the support of The Scientific and Technological Research Council of Turkey (TUBITAK), Project No: 121E088.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.